OpenStudy (anonymous):

inverse of y = -2x + 4

OpenStudy (anonymous):

Basically, you are switching the 'x's and the 'y's, and then isolating to solve for x: x= 3y-4 3y= x+4 y= (x+4)/3 Therefore y^-1 (that's how you write inverse) =(x+4)/3

OpenStudy (anonymous):